OpenStudy (anonymous):

Trouble! lim x->-2 (abs(x^2+2x))/(2x+4)

OpenStudy (anonymous):

sorry \[\lim_{x \rightarrow -2^{-}}\]

OpenStudy (aum):

Factor the numerator and factor the denominator first.

OpenStudy (anonymous):

How would I go about that with the absolute values?

OpenStudy (aum):

\[ \lim_{x \rightarrow -2^{-}} \frac{|x^2+2x|}{2x+4} = \lim_{x \rightarrow -2^{-}} \frac{|x(x+2)|}{2(x+2)} \]

OpenStudy (anonymous):

Alright makes sense

OpenStudy (aum):

When x approaches -2 from the left of -2, what will the numerator be without the absolute bars?

OpenStudy (anonymous):

\[0^{-}\]

OpenStudy (aum):

When x approaches -2 from the left, x < -2. So the x within the absolute bar can be replaced by -x without the absolute bar as both will yield +2 when x->-2. Similarly, (x+2) will be negative when x < -2, so (x+2) within the absolute bars can be replaced by -(x+2) without the absolute bars. \[ \lim_{x \rightarrow -2^{-}} \frac{|x(x+2)}{2(x+2)} = \lim_{x \rightarrow -2^{-}} \frac{-x * -(x+2)}{2(x+2)} = \lim_{x \rightarrow -2^{-}} \frac{-x * -1}{2} = -1 \]

OpenStudy (anonymous):

Ohhh, okay. I was getting confused with the absolute values and limits. I get it now though thanks!

OpenStudy (aum):

You are welcome.